# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v2

@@ III-D1 Synthetic supervised depth loss
since unsupervised depth estimation models based on video sequences cannot recover scale information, in order to enable joint training with synthesized depth which has deterministic depth values, varying with different camera parameters, we adopt the same affine-invariant mean absolute error loss as MiDas[53] and Depth Anything[54]:
Here, and are the scaled and shifted pixels of predicted and synthetic goundtruth .
@@ III-D2 Optimization Strategy
Swami et al. [36] highlighted that multiple networks and optimization losses in unsupervised domain adaptation methods often compete against each other and fail to contribute effectively to the optimization of the depth estimation task. This lack of cooperation among different components makes it challenging to achieve best results. So here, we have set two optimizers to separately train the Depth Net and the Color net, as shown in Fig.2.
Optimizer 1-: Supervised training of the Depth Net on the Syn-Real dataset and the Color Net on the Real world unsupervised depth dataset for predicting color from depth, .
Optimizer 2-: Unsupervised training of the Depth Net on the Real world unsupervised depth dataset and applyed the background textureless regularization loss , .
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.III, Tab.IV, and Tab.I). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.II), perform comparisons at higher resolutions (Tab.V), and benchmark against the lightest and most effective methods (Tab.VIII). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 12, Fig. 15, and Fig. 16).
@@ IV-A Datasets
We utilized two unsupervised monocular depth estimation datasets: KITTI[47] and Cityscapes[9], along with synthetic datasets vKitti2[10], SHIFT[45] and UrbanSyn[46]. KITTI consists of 39,810 training triplets, with 4,424 for validation and 697 for evaluation. It presents diverse scenes with rich color compositions and sporadic non-rigid objects. Cityscapes, tailored for autonomous driving, provides 69,731 monocular triplets from the LeftImg8bit Sequence for training. Primarily urban, it encompasses non-rigid pedestrians and vehicles, often depicted with darker color tones. When using stereo data from the KITTI and Cityscapes datasets, it is important to note that the baseline for KITTI is 54 cm, while for Cityscapes it is 22 cm. vKitti2, an upgraded version of Virtual KITTI, employs the Unity game engine to offer a more photorealistic environment. It provides diverse data types for various computer vision tasks, with 21,260 frames extracted for training. SHIFT, featuring various weather conditions and times, offers dense depth information with a 1 mm resolution and includes simulated pedestrians. We selected 153,000 frames and corresponding synthetic depths from its front-view cameras as our reference synthetic dataset. UrbanSyn[46], designed to address the synth-to-real domain gap, contributes to unprecedented synthetic-only baselines used by domain adaptation (DA) methods. It offers 7,539 images along with depth labels, providing a highly realistic synthetic dataset for various computer vision tasks.
@@ IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.III and Tab.IV with Kitti Eigen Split[47] and Tab.I with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.12 and Fig.15. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[34, 33] with larger params and more complex structures(Tab.VIII).
@@ IV-D Ablation Study
As illustrated in Tab. VI, we conducted experiments on our Back2Color framework by replacing different model structures to demonstrate the effectiveness of our method. While the Depth Net learned to predict dense depth from synthetic data, it also disrupted the model’s latent memory in unsupervised learning and learned to generate sharp boundaries of obstacles. Based on the metrics in Tab. VI, there is an improvement across all metrics, particularly the and metrics representing detail accuracy. At the same time, we also observed that there was no significant difference between using SHIFT and vKitti2. In Tab. VI, we also tested training VADepth from scratch without ImageNet pre-training. The results demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
| Method | Sup. | Size | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | |||
And then, we compared our method with two other noval approaches also joint training with synthetic datasets. Our improvement remains significant in comparison, as shown in Tab.II.
| Method | Sup. | Size | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | |||
| Method | Sup. | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | ||
| PackNet-SfM[23] | m | 0.107 | 0.802 | 4.538 | 0.186 | 0.889 | 0.962 | 0.981 |
| SGDepth[63] | m | 0.107 | 0.768 | 4.468 | 0.186 | 0.891 | 0.963 | 0.982 |
| MonoViT[33] | m | 0.094 | 0.682 | 4.200 | 0.170 | 0.912 | 0.969 | 0.984 |
| PlaneDepth[64] | ms | 0.090 | 0.584 | 4.130 | 0.182 | 0.896 | 0.962 | 0.981 |
| Back2Color | ms | 0.086 | 0.629 | 4.054 | 0.165 | 0.920 | 0.970 | 0.985 |
@@ IV-E Visualization Analysis
The impact of synthetic datasets on real-world unsupervised depth estimation can be analyzed by comparing the joint training with different synthetic datasets, as shown in the Fig.16.
@@
| MonoViT | small | 320x1024 | 33.63 | 159.14 | 0.093 | 0.671 | 4.202 | 0.169 | 0.912 | 0.969 | 0.985 |
| Lite-mono | 8M | 320x1024 | 8.76 | 29.88 | 0.097 | 0.710 | 4.309 | 0.174 | 0.905 | 0.967 | 0.984 |
@@ IV-F1 Why can the network transform information beyond depth
in Fig. 3, Syn-to-Real images exhibit excessive detail surpassing depth information alone, such as tree shadows and lane markings. We attribute this to the Color-Depth-Color process, where depth, predicted from color, retains texture details (such as subtle gradient changes at some boundaries). Consequently, the Color network can recover such gradient changes from the depth map. To validate, we directly trained the network to predict color images from synthesized depth in the synthetic dataset (Fig. 17). The absence of lane markings and shadows in the restored color images confirms our hypothesis.
@@ V Conclusion
In this paper, we presented Back to the Color (Back2Color), a novel method designed to enhance unsupervised depth estimation by leveraging synthetic data. Our method addresses the challenge of aligning synthetic depth information with real-world colors, enabling the transformation of predicted synthetic depth into realistic colors with high fidelity. By joint training on synthetic and real-world datasets using Back2Color, we observed substantial improvements in depth estimation performance. Additionally, we designed a depth network based on the Vision Attention Network (VAN), which allowed our CNN-based network to surpass several Transformer-based depth models. Furthermore, we improved the uncertainty-based loss fusion method by combining the advantages of temporal and spatial reprojection, thoroughly addressing issues related to non-rigid motion. Experimental results demonstrated the effectiveness of our method on both the KITTI and Cityscapes datasets. Overall, Back2Color offers a promising solution for advancing unsupervised depth estimation by effectively integrating synthetic data into the training process.

### v3

@@ III-D1 Synthetic supervised depth loss
Since unsupervised depth estimation models based on video sequences cannot recover scale information, we adopt the affine-invariant mean absolute error loss used by MiDas[55] and Depth Anything[56] for joint training with synthesized depth, which has deterministic values varying with different camera parameters:
where and are the scaled and shifted pixels of predicted and synthetic goundtruth D.
@@ III-D2 Optimization Strategy
Swami et al.[41] highlighted that multiple networks and optimization losses in unsupervised domain adaptation methods often compete against each other and fail to contribute effectively to the optimization of the depth estimation. This lack of cooperation among different components makes it challenging to achieve best results. So here, we have set two optimizers to separately train the Depth Net and the Color net, as shown in Fig.2.
Optimizer 1-: Supervised training of the Depth Net on the Syn-Real samples and the Color Net on the Real world unsupervised depth dataset for predicting color from depth, .
Optimizer 2-: Unsupervised training of the Depth Net on the Real world and applyed the background textureless regularization loss , .
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.I, Tab.II, and Tab.IV). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.VIII). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 13, Fig. 15, and Fig. 16).
@@ IV-A Datasets
We utilized two unsupervised monocular depth estimation datasets: KITTI[26] and Cityscapes[12], along with synthetic datasets vKitti2[13], SHIFT[24], and UrbanSyn[25]. KITTI contains 39,810 training triplets, 4,424 for validation, and 697 for evaluation, featuring diverse scenes with rich colors and occasional non-rigid objects. Cityscapes, designed for autonomous driving, provides 69,731 monocular triplets from the LeftImg8bit Sequence for training, mainly urban scenes with non-rigid pedestrians and vehicles, often in darker tones. The baseline for KITTI stereo data is 54 cm, while for Cityscapes it is 22 cm. vKitti2, an upgraded version of Virtual KITTI, uses the Unity game engine for a more photorealistic environment and provides 21,260 frames for training. SHIFT features various weather conditions and times, offers dense depth information with 1 mm resolution, and includes simulated pedestrians. We selected 153,000 frames and synthetic depths from its front-view cameras. UrbanSyn, designed to address the synth-to-real domain gap, provides 7,539 images with depth labels, offering a highly realistic synthetic dataset for various computer vision tasks.
@@ IV-B Implementation Details
| Method | Sup. | Size | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | |||
| Method | Sup. | Size | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | |||
| DIFFNet[32] | ms | 320x1024 | 0.094 | 0.678 | 4.250 | 0.172 | 0.911 | 0.968 | 0.984 |
| Method | Sup. | Error Metric | Accuracy Metric | |||||
| abs_rel | sq_rel | rmse | rmse_log | a1 | a2 | a3 | ||
| PackNet-SfM[29] | m | 0.107 | 0.802 | 4.538 | 0.186 | 0.889 | 0.962 | 0.981 |
| SGDepth[60] | m | 0.107 | 0.768 | 4.468 | 0.186 | 0.891 | 0.963 | 0.982 |
| MonoViT[2] | m | 0.094 | 0.682 | 4.200 | 0.170 | 0.912 | 0.969 | 0.984 |
| PlaneDepth[61] | ms | 0.090 | 0.584 | 4.130 | 0.182 | 0.896 | 0.962 | 0.981 |
| Back2Color | ms | 0.086 | 0.629 | 4.054 | 0.165 | 0.920 | 0.970 | 0.985 |
@@ IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.I and Tab.II with Kitti Eigen Split[26] and Tab.IV with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.13 and Fig.15. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[1, 2] with larger params and more complex structures(Tab.VIII).
@@ IV-D Ablation Study
As illustrated in Tab. VI, we conducted experiments on our Back2Color framework by replacing different model structures to demonstrate the effectiveness of our method. While the Depth Net learned to predict dense depth from synthetic data, it also disrupted the model’s latent memory in unsupervised learning and learned to generate sharp boundaries of obstacles. Based on the metrics in Tab. VI, there is an improvement across all metrics, particularly the and metrics representing detail accuracy. At the same time, we also observed that there was no significant difference between only using SHIFT and vKitti2. In Tab. VI, we also tested training VADepth from scratch without ImageNet pre-training. The results demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
And then, we compared our method with other noval approaches also joint training with synthetic datasets. Our improvement remains significant in comparison, as shown in Tab.V.
@@ IV-E Visualization Analysis
The impact of synthetic datasets can be analyzed by comparing the joint training with different synthetic datasets, as shown in the Fig.16.
@@
| MonoViT[2] | small | 320x1024 | 33.63 | 159.14 | 0.093 | 0.671 | 4.202 | 0.169 | 0.912 | 0.969 | 0.985 |
| Lite-mono[1] | 8M | 320x1024 | 8.76 | 29.88 | 0.097 | 0.710 | 4.309 | 0.174 | 0.905 | 0.967 | 0.984 |
@@ IV-F1 Why can the network transform information beyond depth
in Fig. 3, Syn-to-Real images exhibit excessive detail surpassing information in depth, such as tree shadows and lane markings. We attribute this to the Color-Depth-Color process, where depth, predicted from color, retains texture details (such as subtle gradient changes at some boundaries). Consequently, the Color Net can recover such gradient changes from the depth map. To validate, we directly trained the network to predict color images from synthesized depth in the synthetic dataset (Fig. 17). The absence of lane markings and shadows in the restored color images confirms our hypothesis.
@@ V Conclusion
In this paper, we presented Back to the Color(Back2Color), a novel framework designed to enhance unsupervised depth estimation by leveraging synthetic data. Our method addresses the challenge of aligning synthetic world with real world colors, enabling the transformation of predicted synthetic depth into realistic colors with high fidelity. By joint training on synthetic and real-world datasets using Back2Color, we observed substantial improvements in depth estimation performance. Additionally, we designed a depth net based on the Vision Attention Network (VAN), which allowed our CNN-based network to surpass several Transformer-based depth models. Furthermore, we improved the uncertainty-based loss fusion method by combining the advantages of temporal and spatial reprojection, thoroughly addressing issues related to non-rigid motion. Experimental results demonstrated the effectiveness of our method on both the KITTI and Cityscapes datasets. Overall, Back2Color offers a promising solution for advancing unsupervised depth estimation by effectively integrating synthetic data into the training process.